NatalieAfton:

Solve the system of equations using the substitution method. y = 7x – 16, y = 2(2x – 5) What is the solution?

AZ:

When we're using the substitution method, we plug in one equation into the other So we have here \(\color{red}{y} = 7x - 16\) \(\color{red}{y} = 2(2x-5)\) we can plug the first equation into the second and we would get 7x - 16 = 2(2x - 5) Can you solve for x?

yourgrandpa:

First, we can substitute in for y in the other equation This will give us Next, we distribute the 2 throughout the parenthesis Now we can solve for x. Begin by adding 16 to each side Now we can subtract 4x from each side Lastly, we can divide each side by 3 Now, we need to substitute in the value that we got for x into one of the equations to find the y value. This becomes So This means that the solution to the system of equations is (2,-2)

NatalieAfton:

thx?

yourgrandpa:

(:

NatalieAfton:

but this needs two answers

yourgrandpa:

Answer 5.0/5 98 pokahontaseals2 Answer is ( 2, -2 ) so x=2 and y=-2 my bad